OpenStudy (anonymous):

The area of triangle formed by the line x+y-8=0 and the pair of straight lines through orgin which passes through the intersection of x^2+y^2-2x-2y-2=0 is?

OpenStudy (klimenkov):

Hope it will help. \(x^2+y^2-2x-2y-2=(x-1)^2+(y-1)^2-4=0\). An equation of the circle.